# AutoRev: Automatic Peer Review System for Academic Research Papers

### v1

@@ 4.4 Grounding Requirements and Inference Efficiency
Table 2: Average token length (Avg. Token Len) and Average passage count (Avg. Passages) for a full paper, our best-configuration: GNN $(5,5)$, and our fastest-configuration: GNN $(5,3)$.
4.5 Fine-Tuning Large Language Models for Review Generation
@@
We select Mistral-Instruct-v0.2 (Mistral) Jiang et al. (2023), inspired by our baseline SEA-EA, which also uses Mistral-Instruct-v0.2 as its underlying backbone model. We also use Llama 3.1 8B (Llama) Dubey et al. (2024), as it is an open-weight state-of-the-art (SOTA) model with strong text generation capabilities for a variety of tasks. We employ LoRA (Low-Rank Adaptation) Hu et al. (2022), for fine-tuning due to its efficiency in adapting large language models to particular tasks while maintaining computational feasibility. This allows us to leverage the pre-trained capabilities of Llama and Mistral while efficiently adapting them for review generation.
We use the trained GNN, as described in Section 4.3, to retrieve relevant passages for each paper that serve as the LLM’s input during fine-tuning. The GPT-4 consolidated reviews for the ICLR 2024 papers, as released by Yu et al. (2024), are regarded as gold standard references for training the LLM to generate final reviews. The dataset split is summarized in Table 3.
@@
Table 3: Number of samples used for training and testing sourced from Yu et al. (2024)
@@
5 Results and Analysis
@@
Table 4: Overall performance on the ICLR-24 reviews test dataset, comparing baselines from Yu et al. (2024) with our fine-tuned models. For both Llama-3.1-8B (Llama) and Mistral-Instruct-v0.2 (Mistral), the configuration with $(5,5)$ achieves the best performance, while $(5,3)$ has the shortest training time. Bold highlights the best scores, and underlined values indicate the second-best. Please note that the GPU configurations for finetuning Llama and Mistral were different. For details, refer to Appendix C.
We evaluate the quality of the generated reviews using ROUGE-1, ROUGE-2, and ROUGE-L (both Recall and F1) Lin (2004), as well as BERTScore-F1 Zhang* et al. (2020), on the test set of GPT-4-consolidated reviews for ICLR 2024 papers. ROUGE measures similarity by analyzing n-grams, with recall indicating how well the generated reviews capture key information from the original papers and the F1 score balancing precision and recall. In contrast, BERTScore assesses semantic similarity in the embedding space, providing a deeper understanding of meaning beyond surface-level text matching.
5.1 Performance Comparison with Baselines
@@
We compare the baselines SEA-E and SEA-EA, two automated review generation approaches, as introduced by Yu et al. (2024). As mentioned in Table 4, our results show that we consistently outperform the baselines in all metrics, with particularly significant improvements in the ROUGE metrics (Recall and F1). Our best-performing configuration, Llama $(5,5)$ (selected based on the highest BERTScore), achieves an improvement of 58.72% over the state-of-the-art (SOTA) baselines across all evaluation metrics. To ensure that these gains are not solely due to using a more capable LLM, we also evaluated our method using Mistral. The results demonstrate that even with Mistral, integrating our GNN-based passage selection leads to substantial improvements, confirming that the performance boost is primarily due to our graph-based retrieval strategy rather than the choice of LLM alone.
5.2 Performance of Our Model
Model Review Section ROUGE-F1 ROUGE-Recall BERTScore-F1 Training Time (hrs)
R-1 R-2 R-L R-1 R-2 R-L
Mistral $(5,3)$ Questions 39.24 9.33 20.00 40.90 9.74 20.91 86.15 4.34
Weaknesses 40.12 9.28 19.56 38.83 8.98 18.98 86.75
Strengths 43.15 11.17 22.07 43.63 11.25 22.47 87.83
Summary 45.18 13.08 25.28 45.35 13.14 25.38 88.09
Llama $(5,3)$ Questions 40.15 9.66 20.02 39.53 9.47 19.85 86.26 3.60
Weaknesses 40.74 9.56 19.63 40.59 9.50 19.60 86.79
Strengths 42.93 11.13 22.09 43.36 11.20 22.44 87.84
Summary 45.65 13.60 25.85 46.67 13.92 26.45 88.23
Mistral $(5,5)$ Questions 39.20 9.00 19.91 40.67 9.32 20.83 86.09 6.44
Weaknesses 40.27 9.21 19.46 39.49 9.01 19.13 86.71
Strengths 43.45 11.49 22.34 44.11 11.61 22.82 87.84
Summary 46.24 14.44 26.32 46.56 14.57 26.48 88.33
Llama $(5,5)$ Questions 39.54 9.31 20.01 38.43 9.02 19.62 86.25 5.30
Weaknesses 40.72 9.66 19.79 40.72 9.59 19.93 86.77
Strengths 43.43 11.76 22.24 43.36 11.69 22.34 87.91
Summary 46.92 14.90 26.74 47.30 14.99 26.92 88.36
Table 5: The sectionwise performance across the ICLR-24 reviews test dataset after finetuning Llama-3.1-8B (Llama) and Mistral-Instruct-v0.2 (Mistral). Per section of the review, the best results have been made bold, while the second-best results have been underlined. Please note that the GPU configurations for finetuning Llama and Mistral were different. For details, refer to Appendix C.
As shown in Table 5, our best-performing configuration, Llama $(5,5)$ consistently outperforms the baselines across ROUGE-F1, ROUGE-Recall, and BERTScore metrics. Llama $(5,5)$ configuration is finetuned with input passages inferred from the GNN $(5,5)$. This highlights its ability to generate reviews that closely align with gold-standard references. The integration of a Graph Attention Network (GAT) enables the model to effectively capture complex dependencies between relevant passages and sentences within the paper, offering the essential context required to generate comprehensive reviews.
Notably, even the fastest variant of our fine-tuned models, Llama $(5,3)$, which requires the shortest training time, performs competitively. Despite using a smaller number of retrieved passages, this variant shows only an average marginal decrease of 0.71% in performance metrics compared to the best-performing configuration, Llama $(5,5)$. Further, it consistently outperforms the state-of-the-art (SOTA) baselines, while reducing training time by 32.07%. This trend is consistently observed across all metrics for Mistral too. It highlights that our GNN-based approach can achieve near-optimal performance with significantly less computational overhead.
As shown in Table 5, a detailed analysis of our model’s performance across different review sections reveals that the Summary and Strengths sections achieve higher scores across ROUGE and BERTScore in our best-performing model, Llama $(5,5)$. This indicates that the model is effective at capturing and utilizing essential context from the paper to generate these key components of the review. Conversely, the Weaknesses and Questions sections receive comparatively lower scores. This indicates that identifying limitations and raising questions require a deeper semantic understanding of the paper as weaknesses may not always be explicitly stated in the passages of the paper. Overall, GNN helps to reflect on the content of the paper comprehensively to generate suitable reviews.
5.3 LLM-as-a-judge Evaluation
Review Review Quality Assessment Criteria
Reviewer Confidence Review Thoroughness Review Constructiveness Review Helpfulness
@@ 5.3 LLM-as-a-judge Evaluation
Table 6: The table reports the qualitative evaluation of model-generated reviews using Gemini-2.0-Flash as an LLM judge. It presents the average scores (out of 5) across four review quality criteria: Reviewer Confidence, Review Thoroughness, Review Constructiveness, and Review Helpfulness, in comparison to gold reviews on the test set. Bold highlights the model best scores, and underlined values indicate the second-best.
As mentioned by Zheng et al. (2023), the LLM-as-a-judge method offers an explainable and affordable way to approximate human preferences, which are otherwise costly to gather. We employ a large language model, Gemini-2.0-Flash [^4] https://deepmind.google/technologies/gemini/ as an automated judge to evaluate the quality of reviews generated by our proposed model as compared to the gold-standard reviews. The evaluation criteria are designed to capture key aspects of review quality and include:
  1. 1.
    Reviewer Confidence: reflects the reviewer’s depth of technical understanding and engagement with the content.
  2. 2.
    Review Thoroughness: measures how well the review covers the methodology, results, and reproducibility.
  3. 3.
    Review Constructiveness: assesses the clarity and usefulness of suggestions for improvement.
  4. 4.
    Review Helpfulness: indicates how well the review identifies important issues and explains them clearly for the authors.
The model is prompted to assign a score from 1 to 5 based on the above criteria. The evaluation prompt is detailed in Appendix D. As shown in Table 6, our best-performing model, Llama $(5,5)$, scores only 9.59% below gold reviews across all criteria. The fastest variant, Llama $(5,3)$, achieves comparable quality with just a 9.78% drop relative to gold reviews while reducing training time by 32.07%. All Llama and Mistral variants consistently score above 3.0, indicating that the reviews generated grasp key contributions and provide relevant feedback with actionable suggestions.
5.4 Ablation Studies
@@
6 Conclusion
In this work, we introduce AutoRev, an automatic peer review generation system designed specifically for academic research papers. Our approach leverages a novel graph-based framework to represent the structure of a paper which enables the effective extraction of important passages using a Graph Neural Network. This strategy of selecting passages and reducing the input length to LLMs addresses the limitations of traditional methods, like finetuning using the entire paper as input Yu et al. (2024), which leads to sub-optimal performance of LLMs due to large input token length Hsieh et al. (2024).
@@ 6 Conclusion
Our experiments on the ICLR 2024 dataset demonstrate that AutoRev consistently outperforms existing baselines, by 58.72% on popular evaluation metrics like ROUGE, and BERTScore.
Overall, AutoRev provides a promising solution for pre-review feedback, aiding both authors and reviewers in the academic publication process. Future work will explore extending this framework to a broader range of academic domains and investigating further improvements in graph construction and neural retrieval techniques to continue advancing the state-of-the-art in automated review generation.
Limitations
While our framework provides a robust and high-performing solution for review generation, we acknowledge that no automatic system can fully replicate human judgment in complex tasks such as peer review, as noted by Ye et al. (2024). Our goal is to assist authors in refining their papers and we aim for our system to support human reviewers. Currently, our system is primarily focused on ICLR papers due to the limited availability of public datasets. However, we believe in the generalization capabilities of our model, given the proven effectiveness of graph-based approaches in various downstream NLP tasks, as discussed in Section 1.
Ethical Considerations
@@ Appendix A Ablation studies
Table 7: Overall performance of Mistral across different $(k,m)$ settings. Metrics include ROUGE-F1 (R-1, R-2, R-L), ROUGE-Recall (R-1, R-2, R-L), BERTScore F1 and training time (in hours). Bold highlights the best scores, and underlined values indicate the second-best.
@@
Table 8: Overall performance of Llama across different $(k,m)$ settings. Metrics include ROUGE-F1 (R-1, R-2, R-L), ROUGE-Recall (R-1, R-2, R-L), BERTScore F1 and training time (in hours). Bold highlights the best scores, and underlined values indicate the second-best.

### v2

@@ 4.4. Grounding Requirements and Inference Efficiency
4.5. Fine-Tuning Large Language Models for Review Generation
@@
We select Mistral-Instruct-v0.2 (Mistral) [^7] https://arxiv.org/abs/2310.06825, inspired by our baseline SEA-EA, which also uses Mistral-Instruct-v0.2 as its underlying backbone model. We also use Llama 3.1 8B (Llama) [^8] https://arxiv.org/abs/2407.21783, as it is an open-weight state-of-the-art (SOTA) model with strong text generation capabilities for a variety of tasks. We employ LoRA (Low-Rank Adaptation) (Hu et al., 2022), for fine-tuning due to its efficiency in adapting large language models to particular tasks while maintaining computational feasibility. This allows us to leverage the pre-trained capabilities of Llama and Mistral while efficiently adapting them for review generation.
We use the trained GNN, as described in Section 4.3, to retrieve relevant passages for each paper that serve as the LLM’s input during fine-tuning. The human-written peer reviews for ICLR 2024 papers, which are consolidated with GPT-4, served as the ground truth for fine-tuning the LLM used to generate the final reviews. These reviews are released as a part of the dataset released by Yu et al. (2024). The dataset split is summarized in Table 3.
Table 3. Number of samples used for training and testing sourced from Yu et al. (2024)
@@
5. Results and Analysis
Table 4. Overall performance on the ICLR-24 reviews test dataset, comparing M-7B, SEA-E, SEA-EA as baselines from Yu et al. (2024) with our fine-tuned models. For both Llama-3.1-8B (Llama) and Mistral-Instruct-v0.2 (Mistral), the configuration with $(5,5)$ achieves the best performance, while $(5,3)$ has the shortest training time. Bold highlights the best scores, and underlined values indicate the second-best. Please note that the GPU configurations for finetuning Llama and Mistral were different. For details, refer to Appendix C.
@@ 5. Results and Analysis
This section details our system’s performance analysis. Section 5.1 reports the evaluation of generated review quality using ROUGE-1, ROUGE-2, and ROUGE-L (Recall and F1) (Lin, 2004), as well as BERTScore-F1 (Zhang* et al., 2020), on the GPT-4-consolidated reviews test set for ICLR 2024 papers. ROUGE measures n-gram overlap, with recall reflecting content coverage and F1 balancing precision and recall, while BERTScore captures semantic similarity in embedding space. Section 5.2 further employs the Aspect Coverage (ACov) metric (Yuan et al., 2022), designed specifically for assessing peer-review quality.
Given the limitations of automated metrics, we also conduct human evaluations (Section 5.3) to assess review acceptability. To address the inherent scalability challenges of manual assessment, we employ an LLM-as-a-Judge framework (Section 5.4), which shows strong correlation with human judgments.
5.1. Performance Comparison with Baselines
@@
We compare the performance of our models against baselines sourced from Yu et al. (2024). As reported in Table 4, the M-7B baseline performs direct inference with Mistral guided by an instruction prompt to generate peer reviews, SEA-E baseline leverages standardized fine-tuning data to produce more constructive reviews, while the SEA-EA baseline further introduces a mismatch score to evaluate the consistency between paper content and generated reviews. Our results suggest that we consistently outperform the baselines on all metrics, with significant improvements in the ROUGE metrics (Recall and F1). Our best-performing configuration, Llama $(5,5)$ (selected based on the highest BERTScore), achieves an improvement of 58.72% over the state-of-the-art (SOTA) baselines across all evaluation metrics. To ensure that these gains are not solely due to using a more capable LLM, we also evaluate our method using Mistral. The results demonstrate that even with Mistral, integrating our Multi Modal GNN-based passage selection method leads to substantial improvements, confirming that the performance boost is primarily due to our graph-based retrieval strategy rather than the choice of the LLM alone. The integration of a Graph Attention Network (GAT) enables the model to effectively capture complex dependencies between relevant passages and sentences within the paper, offering the essential context required to generate comprehensive reviews.
Our fastest, second-best, fine-tuned model, Llama $(5,3)$, performs competitively while being highly efficient. It reduces the training time by 32.07% with only a 0.71% drop in performance compared to our best model, consistently outperforming all baselines. This shows our GNN-based approach achieves near-optimal performance with significantly less computational overhead.
5.2. Aspect Coverage Analysis
As discussed by Yuan et al. (2022), each comment in a peer review typically addresses diverse aspects of an academic research paper, such as Summary (SUM), Motivation/Impact (MOT), Originality (ORI), Soundness/Correctness (SOU), Substance (SUB), Replicability (REP), Meaningful Comparison (CMP), and Clarity (CLA). These cannot be reliably assessed by conventional automatic metrics like ROUGE-F1 or BERTScore. To address this, Yuan et al. (2022) introduces an Aspect Tagger which can be used to compute Aspect Coverage (ACov) that measures the proportion of aspects covered by a peer-review. We evaluate the two best-performing models Llama $(5,5)$ and Llama $(5,3)$ using this tagger.
Table 5. Aspect Coverage (%) of model-generated reviews evaluated using the Aspect Tagger (Yuan et al., 2022). Summary (SUM), Clarity (CLA), Soundness/Correctness (SOU), Substance (SUB), Originality (ORI), Meaningful Comparison (CMP), Replicability (REP), and Motivation/Impact (MOT).
Model SUM CLA SOU SUB ORI CMP REP MOT
Llama (5,3) 100.0 99.6 98.9 95.2 92.6 79.0 72.0 66.8
Llama (5,5) 100.0 99.3 99.3 95.9 90.0 84.5 71.6 66.8
As shown in Table 5, both variants achieve strong coverage for Summary, Clarity, and Soundness, indicating the reliability of our model-generated reviews in capturing these core aspects. However, the aspect coverage for Motivation/Impact and Replicability is comparatively lower. Overall, the results suggest that our model captures most key aspects, indicating that the generated reviews are comprehensive and address the quality of multiple dimensions of a research paper.
5.3. Human Evaluation
Understanding the limitations of automatic evaluation metrics as highlighted by Hanna and Bojar (2021) and Schluter (2017), we conduct human evaluations to assess the acceptability of the generated reviews. We model our evaluation around two distinct aspects of review quality. The first is a Qualitative Assessment, where the generated review is evaluated in isolation across four key criteria: Reviewer Confidence, Review Thoroughness, Review Constructiveness, and Review Helpfulness. The second is a Factual Validity Assessment, which ensures the review’s specific claims are factually grounded in the source document. Since human evaluation is very costly, we conduct each of the two tasks on a subset of the test dataset, as described in their individual subsections. Additionally, we only considered Llama $(5,3)$ and Llama $(5,5)$, our best performing models for the evaluation process.
5.3.1. Qualitative Assessment
A total of 30 papers are chosen at random from our test set. Each of these papers are associated with 3 reviews, one Gold, one from Llama $(5,3)$, and one from Llama $(5,5)$, thus, resulting in a total set of 90 reviews. A total of 27 annotators are chosen, and each annotator is given 10 reviews to annotate. The evaluation criteria is designed to capture key aspects of review quality and include:
  1. (1)
    Reviewer Confidence (RCf): Reflects the reviews’ depth of technical understanding and engagement with the content.
  2. (2)
    Review Thoroughness (RT): Measures how well the review covers the methodology, results, and reproducibility.
  3. (3)
    Review Constructiveness (RCt): Assesses the clarity and usefulness of suggestions for improvement.
  4. (4)
    Review Helpfulness (RH): Indicates how well the review identifies important issues and explains them clearly to the authors.
Each annotator is tasked to assign a score from 1 to 5 for each of the above criteria. The evaluation guidelines are detailed in Appendix D. We end up with each review being annotated by 3 annotators. Score for a review is averaged across all its annotators, and the final score of a model, is the average score across all its reviews. The results of this experiment are summarized in Table 6.
Table 6. Results of our human qualitative assessment. This table shows average scores (on a scale of 1-5) comparing gold-standard reviews with those generated by our two best-performing Llama models across four key criteria: Reviewer Confidence (RCf), Review Thoroughness (RT), Review Constructiveness (RCt), and Review Helpfulness (RH).
Review RCf RT RCt RH
Gold 3.54 3.49 3.26 3.41
Llama (5,3) 3.37 3.30 3.27 3.34
Llama (5,5) 3.27 3.24 3.19 3.30
5.3.2. Factual Validity Assessment
A total of 10 papers are chosen at random from our test set. Each of these papers are associated with 3 reviews, one Gold, and two generated by our best performing models. Thus, the final evaluation set consists of 10 papers, each associated with 3 reviews, one Gold, one from Llama $(5,3)$, and one from Llama $(5,5)$. These 10 papers are circulated among 10 annotators. It is ensured that a single paper is annotated by at least 2 annotators. Annotators are instructed to evaluate each statement within a review based on its factual grounding in the full source paper. For each point, they assign a rating from a four-point scale: ‘Completely Agree’, ‘Mostly Agree’, ‘Mostly Disagree’, or ‘Completely Disagree’. The evaluation guidelines are detailed in Appendix D. The results of this experiment are summarized in Table 7 and Table 8.
Table 7. Detailed results of the Factual Validity Assessment. This table presents the percentage distribution of human agreement ratings for claims made in reviews from our Llama models and the gold standard. The abbreviations are: Completely Agree (CA), Mostly Agree (MA), Mostly Disagree (MD), and Completely Disagree (CD).
Review CA MA MD CD
Gold 39.12% 37.85% 16.40% 6.62%
Llama (5,3) 32.64% 31.94% 22.92% 12.50%
Llama (5,5) 32.27% 32.98% 23.05% 11.70%
Table 8. Detailed results of the Factual Validity Assessment. This table shows the overall Positive Rating Percentage, calculated by summing the “Completely Agree” and “Mostly Agree” categories from Table 7, which indicates the factual soundness of the reviews.
Review Positive Rating %
Gold 76.97%
Llama (5,3) 65.25%
Llama (5,5) 64.58%
5.4. LLM-as-a-Judge Evaluation
As mentioned by Zheng et al. (2023), the LLM-as-a-Judge method offers an explainable and affordable way to approximate human preferences, which are otherwise costly to gather. We employ a large language model, Gemini-2.0-Flash [^9] https://deepmind.google/technologies/gemini/ as an automated judge to evaluate the quality of reviews generated by our proposed model as compared to the gold-standard reviews. For the LLM-as-a-judge task, the evaluation criteria and prompting guidelines are adapted from those used in our human qualitative assessment, which are detailed in Section 5.3.1. Since LLM-as-a-Judge evaluation is scalable and much cheaper to undertake, we evaluate all our models against the gold review, over the entire test dataset. A key finding from our evaluation is the strong correlation between the LLM-as-a-Judge scores (Table 9) and the results from our human annotators (Table 6). This alignment suggests that the LLM-as-a-Judge is a reliable and valid proxy for assessing review quality in this domain.
Table 9. Qualitative evaluation of model-generated reviews using Gemini-2.0-Flash as the LLM judge. The table presents the average scores (out of 5) for all model configurations and the gold standard across four criteria: Reviewer Confidence (RCf), Review Thoroughness (RT), Review Constructiveness (RCt), and Review Helpfulness (RH). Bold values indicate the best score in each column (excluding Gold), and underlined values indicate the second-best.
Review RCf RT RCt RH
@@ 5.4. LLM-as-a-Judge Evaluation
5.5. Ablation Studies
@@
6. Limitations
While our framework provides a robust and high-performing solution for review generation, we acknowledge that no automatic system can fully replicate human judgment in complex tasks such as peer reviewing, as noted by Ye et al. (2024). Our goal is to assist authors in refining their papers and we aim for our system to support human reviewers. Currently, our system is primarily focused on ICLR 2024 papers due to the limited availability of public datasets containing consolidated reviews. However, we believe in the generalization capabilities of our model, given the proven effectiveness of graph-based approaches in various downstream NLP tasks, as discussed in Section 1.
Our reliance on human written reviews, consolidated using GPT-4 represents another limitation. While multi-review datasets more closely reflect the diversity of real peer review, they are often noisy and contradictory, which can hinder model learning. Consolidated reviews provide a clearer and more consistent supervision signal, but at the cost of reducing the richness of differing reviewer perspectives. Future work should explore hybrid approaches that retain both consistency and diversity.
7. Ethical Considerations
8. Conclusion
In this work, we introduce AutoRev, an automatic peer review generation system designed specifically for academic research papers. Our approach leverages a novel graph-based framework to represent the structure of a paper which enables the effective extraction of important passages using a Graph Neural Network. This strategy of selecting passages and reducing the input length to LLMs addresses the limitations of traditional methods, like finetuning using the entire paper as input (Yu et al., 2024), which leads to sub-optimal performance of LLMs due to large input token length (Hsieh et al., 2024).
@@ 8. Conclusion
Our experiments on the ICLR 2024 dataset demonstrate that AutoRev consistently outperforms existing baselines, by 58.72% on popular evaluation metrics like ROUGE, and BERTScore. These findings are further strongly backed by human evaluations, where the generated reviews were rated nearly as constructive and acceptable as human-written ones, suggesting AutoRev could meaningfully reduce reviewer workload.
Beyond these performance gains, we emphasize AutoRev’s broader role in the web ecosystem. Academic research papers represent some of the most authoritative and trusted content on the web. By accelerating the peer-review workflow and enabling scalable, automated feedback, AutoRev contributes to the timely dissemination of credible scholarly knowledge online. In doing so, it supports not only authors and reviewers but also the wider web community that depends on accurate, trustworthy research for discovery, indexing, and downstream applications. Future work will explore extending this framework to a broader range of academic domains and investigating further improvements in graph construction and neural retrieval techniques to continue advancing the state-of-the-art in automated review generation.
@@ Appendix A Ablation studies
Table 10. Overall performance of Mistral across different $(k,m)$ settings. Metrics include ROUGE-F1 (R-1, R-2, R-L), ROUGE-Recall (R-1, R-2, R-L), BERTScore F1 and training time (in hours). Bold highlights the best scores, and underlined values indicate the second-best.
@@
Table 11. Overall performance of Llama across different $(k,m)$ settings. Metrics include ROUGE-F1 (R-1, R-2, R-L), ROUGE-Recall (R-1, R-2, R-L), BERTScore F1 and training time (in hours). Bold highlights the best scores, and underlined values indicate the second-best.